Circular No. 7197

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 L1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet that he discovered in images taken with the C2 and C3 coronagraphs aboard SOHO. No tail was evident. Astrometric measurements made by Biesecker (and reduced by G. V. Williams), together with orbital elements by B. G. Marsden, are given on MPEC 1999-L23.

1999 UT  α_{2000}  δ_{2000}
June 5.013  4^{h}51.^{m}1  +19^{o}40′

XTE J1550–564 M. Buxton, Research School of Astronomy and Astrophysics, Aus-

tralian National University; and S. Vennes, D. Wickramasinghe, and L. Ferrario, Astrophysical Theory Centre, Australian National University, write: “We have obtained optical spectroscopy of the blackhole candidate XTE J1550–564 during 1999 using the Anglo-Australian Telescope on Apr. 13 and 14, and the 2.3-m telescope at Siding Spring on Mar. 11, 24–27, and May 11. We have measured the equivalent width of Hα with the following results: Mar. 11.67–11.75 UT, EW = −1.72 ± 0.02 nm; 24.59–24.76, −1.56 ± 0.09 nm; 25.57–25.68, −1.38 ± 0.03 nm; 26.59–26.63, −1.84 ± 0.20 nm; 27.61–27.65, −1.91 ± 0.10 nm; Apr. 13.50–13.69, −2.41 ± 0.09 nm; 14.49–14.72, −2.48 ± 0.18 nm; May 11.48–11.70, −3.49 ± 0.07 nm. We clearly detect Hα emission in all spectra, together with He I 587.5-nm and sometimes the Paschen series, but no Hβ, He II, or Bowen blend is evident. The interstellar lines at Na D and 628.4 nm are also present. These spectra contradict the result of Soria and Wu (IAUC 7184), in that our data, which bracket their spectrum taken at the AAT on Apr. 10, show Hα in emission and not absorption; therefore we do not confirm detection of the secondary. We also find that the continuum steadily decreases over time and, as a consequence, the EW of Hα increases. The decreasing continuum corresponds to the decrease of optical emission as reported by Jain et al. (IAUC 7114, 7187).”

V382 VELORUM Corrigendum. On IAUC 7196, line 10, for 1.9 × 10^{33} read 4.8 × 10^{33}

1999 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green